# One Look is Enough: A Novel Seamless Patchwise Refinement for Zero-Shot Monocular Depth Estimation Models on High-Resolution Images

### v1

@@ Abstract
Zero-shot depth estimation (DE) models exhibit strong generalization performance as they are trained on large-scale datasets. However, existing models struggle with high-resolution images due to the discrepancy in image resolutions of training (with smaller resolutions) and inference (for high resolutions). Processing them at full resolution leads to decreased estimation accuracy on depth with tremendous memory consumption, while downsampling to the training resolution results in blurred edges in the estimated depth images. Prevailing high-resolution depth estimation methods adopt a patch-based approach, which introduces depth discontinuity issues when reassembling the estimated depth patches and results in test-time inefficiency. Additionally, to obtain fine-grained depth details, these methods rely on synthetic datasets due to the real-world sparse ground truth depth, leading to poor generalizability. To tackle these limitations, we propose Patch Refine Once (PRO), an efficient and generalizable tile-based framework. Our PRO consists of two key components: (i) Grouped Patch Consistency Training that enhances test-time efficiency while mitigating the depth discontinuity problem by jointly processing four overlapping patches and enforcing a consistency loss on their overlapping regions within a single backpropagation step, and (ii) Bias Free Masking that prevents the DE models from overfitting to dataset-specific biases, enabling better generalization to real-world datasets even after training on synthetic data. Zero-shot evaluation on Booster, ETH3D, Middlebury 2014, and NuScenes demonstrates into which our PRO can be well harmonized, making their DE capabilities still effective for the grid input of high-resolution images with little depth discontinuities at the grid boundaries. Our PRO runs fast at inference time.
@@
Figure 1: Qualitative comparison of patch-based depth estimation models (PatchFusion [23], PatchRefiner [24] and PRO (Ours)) for high-resolution images. PatchRefiner [24] and PRO (ours) also visualize the residuals to effectively highlight depth discontinuity issues in the 4th and 6th columns. All models performed depth estimations on $4\times 4$ patches of the 1st and 3rd input images. Zoom-in can help distinguish the grid lines (depth discontinuities) of the $4\times 4$ patches. Our proposed PRO achieves smooth boundary transitions without depth discontinuity artifacts along the grid boundaries.
@@ 1 Introduction
Monocular Depth Estimation (MDE) [10, 33, 36, 27, 43, 13, 14] has been widely investigated as the demand for 3D information continues to grow in autonomous driving, robotics, and virtual reality applications. After Midas [30] proposed an MDE network trainable with mixed training dataset, numerous zero-shot MDE networks [31, 9, 2, 46, 47, 19, 11] have been studied to predict depth for unseen real-world images. However, due to the architectural limitation or limited image resolutions of training dataset, most zero-shot MDE networks are trained specifically with the images of low resolutions (e.g. 384$\times$384, 518$\times$518). Consequently, when high resolution images are inputted to these models, the resulting estimated depth images tend to contain disrupted overall structures with low-frequency artifacts, although yielding improved edge details [26, 7].
@@
Patch-based methods [26, 24, 23] have shown promising results by splitting high-resolution images into patches for DE, mitigating memory consumption issues while achieving remarkable performance. However, they suffer from depth discontinuity problems (e.g. boundary artifacts along grid boundaries) which occur when independent DE patches are reassembled to construct a complete (whole) depth map because depth continuity is maintained within each patch but not between patches. Previous methods [24, 23] alleviate this depth discontinuity problem (i) by incorporating a consistency loss during training and (ii) by ensemble averaging at test time which slows down inference speeds, making it impractical for real-world applications.
@@
Figure 2: Performance and efficiency comparison on Middlebury 2014 [34]. The area of each circle represents the inference time. Circles of the same color represent the same model with different patch numbers ($\mathrm{P}=16$ for small-sized yellow and grey circles, $\mathrm{P}=177$ for large-sized yellow and grey circles). Our model (PRO) achieves the best performance in terms of edge errors ($\mathrm{D}^{3}\mathrm{R}$) and depth errors (AbsRel) while maintaining the fastest inference time. Inference time for all models is measured on the RTX 4090.
@@
A common strategy for training high-resolution depth estimation is to train models on real-world high-resolution datasets [17, 20, 39, 42, 34] or leverage high-resolution synthetic datasets [38, 41, 40, 18]. Real-world datasets have the advantage of a smaller domain gap compared to synthetic datasets, but their ground truth (GT) depths are often sparse, particularly around edges, where supervision is crucial. Additionally, acquiring real-world dense depth data is challenging, limiting the DE learning for high-resolution images. On the other hand, synthetic datasets provide dense GT depth with fine details, but models trained on synthetic datasets often struggle with zero-shot inference on real-world data due to the domain gap [21]. Moreover, we observe that transparent objects do not have accurate depth annotations in the UnrealStereo4K dataset [38] that is the only synthetic dataset with 4K depth labeling, as they are annotated with the depth of the background behind the transparent objects.
@@
To address the aforementioned problems, we propose Patch Refine Once (PRO), a novel refinement model that enables DE models to generate accurate DE results via only a single refinement per patch during inference (i.e.,“One Look”). Our PRO consists of two strategies: Grouped Patch Consistency Training and Bias Free Masking. Fig. 1 shows qualitative comparison of patch-based depth estimation models for high-resolution images. As shown, our PRO yields depth refinement results with little depth discontinuities along the grid boundaries while the very recent two state-of-the-art (SOTA) models exhibit depth discontinuity artifacts. Fig. 2 illustrates performance and efficiency comparison for SOTA depth refinement models and our PRO. As observed, our PRO provides the best performance in terms of $\mathrm{D}^{3}\mathrm{R}$, AbsRel and inference speed. Our contribution is summarized as:
@@
  • We propose Grouped Patch Consistency Training (GPCT) strategy that can be well harmonized with existing DE models to mitigate the depth discontinuity problem for patch-wise DE approaches on high-resolution images. The GPCT does not require test-time ensemble, allowing a 12$\times$ faster inference speed compared to the SOTA patch-based method;
@@
  • We introduce Bias Free Masking (BFM) to selectively apply supervision signals by masking out unreliable regions in images. The unreliable regions are the regions that have inaccurately annotated depth in GT, which is often observed in the UnrealStereo4K dataset. This is problematic without our BFM that efficiently prevents depth refinement models from overfitting to domain specific bias;
@@
  • Our PRO achieves SOTA performance in zero-shot evaluation on Booster [29], ETH3D [35], Middlebury 2014 [34], and NuScenes [4] across all metrics, and runs efficiently at inference time (12$\times$ faster than PatchRefiner with 177 patches), compared to the very recent depth refinement methods for high-resolution images.
@@ 2.1 Zero-shot Monocular Depth Estimation
Early works on MDE models [1, 25, 10, 33, 49] focused on training on each dataset [12, 37], resulting in high performance on the trained dataset with degenerate performance on unseen datasets because of domain gap between datasets. Zero-shot MDE models that perform well on unseen (in-the-wild) images have been widely studied to improve generalization. MegaDepth [22] and DiverseDepth [48] construct large-sized datasets by collecting images from the Internet, improving adaptability on images from diverse scenes and conditions. They encourage efforts to scale up datasets to enhance zero-shot performance.
MiDaS [30] proposes a scale and shift invariant loss, which enables MDE networks to be trained on mixed diverse dataset, making the model robust to unseen datasets. DPT [31], Omnidata [9], and MiDaS v3.1 [2] improve DE performance by applying transformer architectures to replace CNN architectures. DepthAnything [46] introduces semi-supervised learning that utilizes 62M unlabeled images as pseudo labels. A dramatic increase in dataset size has demonstrated robust generalization performance. DepthAnythingv2 [47] points out that real-world depth datasets contain label noise and lack fine details. It is trained with pseudo GT labels obtained from a teacher model trained on a synthetic dataset, achieving high-quality depth maps.
Instead of scaling up datasets, some studies deploy prior knowledge from Diffusion Models for DE. Marigold [19] fine-tunes a pretrained Stable Diffusion [32] model on a relatively small synthetic dataset, showing competitive results. Geowizard [11] jointly estimates depth and surface normals by utilizing self-attention across different domains to enhance geometric consistency. However, zero-shot MDE models are trained on datasets of low resolutions (e.g., 384$\times$384, 518$\times$518), which results in degenerate performance on high resolution images. When high-resolution images are downsampled to the training resolution for inference, fine details are lost. On the other hand, performing inference with a larger resolution input preserves details but degrades depth accuracy.
@@ 2.2 High-Resolution Depth Estimation
High-resolution DE models aim to predict accurate depth while preserving fine details. SMD-Net [38] utilizes an implicit function [5] to predict mixture density for precise DE at object boundaries. Dai et al. [7] introduce a Poisson fusion-based depth map optimization method by applying guided filtering in a self-supervised learning framework. SDDR [21] proposes a self-distilled depth refinement method that generates pseudo edge labels, addressing local consistency issues and edge deformation noise with Noisy Poisson Fusion. Patch-based high-resolution DE methods select patches and merge patch-wise results to improve depth details. BoostingDepth [26] introduces a patch selection process based on edge density, and employs iterative refinement through multi-resolution depth merging. PatchFusion [23] eliminates the need for patch selection by splitting patches according to a predefined method and utilizes shifted window self-attention to fuse global and local information. PatchRefiner [24] initializes depth estimation with a low-resolution depth map and predicts residuals, and proposes a detail and disentangling (DSD) loss for training on both real-world and synthetic datasets. However, patch-based methods inherently suffer from depth discontinuities at patch boundaries, as they emphasize local information over global consistency, or maintain depth continuity only within individual patches. To address this problem, prior works [23, 24] employ a significantly large number of patches for test-time ensemble, which considerably slows down inference speed. By contrast, our PRO (Patch Refine Once) model achieves fine-grained details efficiently by performing only a single refinement per patch during inference, despite utilizing a tile-based approach.
2.3 Training Data for High-resolution Depth Estimation
High-resolution DE requires datasets with fine-grained depth details and minimal label noise. Real-world datasets [12, 37, 4, 15, 22] are often sparse due to the limitations of LiDAR or contain inaccurate depth boundaries when obtained through Structure-from-Motion (SfM). Therefore, previous researches [7, 26] have trained the networks using a small number of relatively dense real-world datasets [44, 20, 34]. Some studies [7, 21] utilize pseudo labels for self-supervision, but these labels often contain noise. Other works [23, 24] leverage synthetic datasets for more accurate depth, resulting in degenerate generalization abilities due to the domain gaps. To take advantage of dense depth annotations in synthetic datasets without compromising generalization performance, we introduce Bias Free Masking (BFM), which selectively identifies the regions where supervision is applied. It mitigates overfitting to synthetic dataset biases and enhances robustness in zero-shot high-resolution DE.
@@ 3 Method
Figure 3: Labeling error examples in UnrealStereo4K [38] dataset.
@@
We describe our Patch Refine Once (PRO) model. In Sec. 3.1, we explain the process of PRO. Then, Grouped Patch Consistency Training (GPCT) and Bias Free Masking (BFM), which are the key contributions of our paper, are described in Sec. 3.2 and Sec. 3.3, respectively.
@@ 3.1 Overview of Patch Refine Once (PRO)
To complement the local details, we take the following steps. First, we crop the original image $\mathbf{I}$ into $N$ grid-based patches $\{\mathbf{P}^{i}\}_{i=1}^{N}$. Each individual patch $\mathbf{P}^{i}$ is then resized and fed into $\Psi$ to estimate the corresponding fine depth map $\mathbf{D}_{\mathrm{f}}^{i}$. We extract two sets of five-level features from the decoder of $\Psi$: (i) $\mathbf{F}_{\mathrm{c}}=\{\mathbf{f}_{\mathrm{c,j}}^{i}\}_{j=1}^{5}$ for the resized image and (ii) $\mathbf{F}_{\mathrm{f}}=\{\mathbf{f}_{\mathrm{f},j}^{i}\}_{j=1}^{5}$ for $\mathbf{P}^{i}$. To align $\mathbf{F}_{\mathrm{c}}$ with $\mathbf{F}_{\mathrm{f}}$, we apply the $\mathsf{ROI}$ operation [16] to extract patch-aligned features from $\mathbf{F}_{\mathrm{c}}$. The patch-level features $\mathbf{F}_{\mathrm{f}}^{i}$ and the corresponding $\mathsf{ROI}$-extracted features are subsequently passed into a fusion module within the residual prediction network $\theta$. The fusion module integrates these features using a wavelet transform [8] and shallow convolutional layers, effectively combining global and local information. Finally, the refined depth map for $\mathbf{P}^{i}$, denoted as $\mathbf{D}_{\mathrm{refine}}^{i}$, is obtained as:
$\mathbf{D}_{\mathrm{refine}}^{i}=\theta(\mathsf{concat}(\mathbf{P}^{i},\mathsf{ROI}(\mathbf{D}_{\mathrm{c}}),\mathbf{D}_{\mathrm{f}}^{i})),$ (1)
where $\mathsf{ROI}(\mathbf{D}_{\mathrm{c}})=\mathbf{D}_{\mathrm{c}}^{i}$ is a region of interest extraction operator for $\mathbf{D}_{\mathrm{c}}$ for $\mathbf{P}^{i}$, yielding $\mathbf{D}_{\mathrm{c}}^{i}$, and $\mathsf{concat}$ denotes channel-wise concatenation. Unlike previous methods [23, 24] where separate networks are trained to predict $\mathbf{D}_{\mathrm{c}}$ and $\mathbf{D}_{\mathrm{f}}$ individually, we utilize the same pretrained zero-shot MDE network $\Psi$ and train only the residual prediction network as shown in Fig. 4-(a).

### v2

@@ Abstract
Zero-shot depth estimation (DE) models exhibit strong generalization performance as they are trained on large-scale datasets. However, existing models struggle with high-resolution images due to the discrepancy in image resolutions of training (with smaller resolutions) and inference (for high resolutions). Processing them at full resolution leads to decreased estimation accuracy on depth with tremendous memory consumption, while downsampling to the training resolution results in blurred edges in the estimated depth images. Prevailing high-resolution depth estimation methods adopt a patch-based approach, which introduces depth discontinuity issues when reassembling the estimated depth patches, resulting in test-time inefficiency. Additionally, to obtain fine-grained depth details, these methods rely on synthetic datasets due to the real-world sparse ground truth depth, leading to poor generalizability. To tackle these limitations, we propose Patch Refine Once (PRO), an efficient and generalizable tile-based framework. Our PRO consists of two key components: (i) Grouped Patch Consistency Training that enhances test-time efficiency while mitigating the depth discontinuity problem by jointly processing four overlapping patches and enforcing a consistency loss on their overlapping regions within a single backpropagation step, and (ii) Bias Free Masking that prevents the DE models from overfitting to dataset-specific biases, enabling better generalization to real-world datasets even after training on synthetic data. Zero-shot evaluations on Booster, ETH3D, Middlebury 2014, and NuScenes demonstrate that our PRO can be seamlessly integrated into existing depth estimation models. It preserves the performance of original depth estimation models even under grid-based inference on high-resolution images, exhibiting minimal depth discontinuities along patch boundaries. Moreover, our PRO achieves significantly faster inference speed compared to prior patch-based methods.
@@
Figure 1: Qualitative comparison of patch-based depth estimation models (PatchFusion [23], PatchRefiner [24] and PRO (Ours)) for high-resolution images. PatchRefiner [24] and PRO (Ours) also visualize the residuals to effectively highlight depth discontinuity issues in the 4th and 6th columns. All models perform depth estimations on $4\times 4$ patches of input images. Zoom-in can help distinguish the grid lines (depth discontinuities) of the $4\times 4$ patches. Our proposed PRO achieves smooth boundary transitions without depth discontinuity artifacts along the grid boundaries.
@@ 1 Introduction
Monocular Depth Estimation (MDE) [10, 33, 36, 27, 43, 13, 14] has been widely investigated as the demand for 3D information continues to grow in autonomous driving, robotics, and virtual reality applications. After Midas [30] proposed an MDE network trainable with mixed training datasets, numerous zero-shot MDE networks [31, 9, 2, 46, 47, 19, 11] have been studied to predict depth for unseen real-world images. However, due to the architectural limitation or limited image resolutions of training datasets, most zero-shot MDE networks are trained specifically with the images of low resolutions (e.g. 384$\times$384, 518$\times$518). Consequently, when high resolution images are fed into these models, the resulting estimated depth images tend to contain disrupted overall structures with low-frequency artifacts, although yielding improved edge details [26, 7].
@@
Patch-based methods [26, 24, 23] have shown promising results by splitting high-resolution images into patches for DE, mitigating memory consumption issues while achieving remarkable performance. However, they suffer from depth discontinuity problems (e.g. boundary artifacts along grid boundaries) which occur when independent DE patches are reassembled to construct a complete (whole) depth map because depth continuity is maintained within each patch but not between patches. Previous methods [24, 23] alleviate this depth discontinuity problem (i) by incorporating a consistency loss during training and (ii) by ensemble averaging at test time which slows down inference speed, making it impractical for real-world applications.
@@
Figure 2: Comparison of performance and efficiency on Middlebury 2014 [34]. The area of each circle represents the inference time. Circles of the same color represent the same model with different patch numbers ($\mathrm{P}=16$ for small-sized yellow and grey circles, $\mathrm{P}=177$ for large-sized yellow and grey circles). Our model (PRO) achieves the best performance in terms of edge errors ($\mathrm{D}^{3}\mathrm{R}$) and depth errors (AbsRel) while maintaining the fastest inference time. Inference time for all models is measured on the RTX 4090.
@@
A common strategy for training high-resolution depth estimation is to train models on real-world high-resolution datasets [17, 20, 39, 42, 34] or leverage high-resolution synthetic datasets [38, 41, 40, 18]. Real-world datasets have the advantage of a smaller domain gap compared to synthetic datasets, but their ground truth (GT) depths are often sparse, particularly around edges, where supervision is crucial. Additionally, the difficulty of acquiring dense depth data in real-world scenarios hinders the training of DE models for high-resolution images. On the other hand, synthetic datasets provide dense GT depth with fine details, but models trained on synthetic datasets often struggle with zero-shot inference on real-world data due to the domain gap [21]. Moreover, we observe that transparent objects are inaccurately labeled in the UnrealStereo4K dataset [38], the only synthetic dataset providing 4K depth annotations, as their depth values correspond to the background behind them.
@@
To address the aforementioned problems, we propose Patch Refine Once (PRO), a novel refinement model that enables DE models to generate accurate depth predictions with only a single refinement per patch at inference time (i.e.,“One Look”). Our PRO consists of two strategies: Grouped Patch Consistency Training and Bias Free Masking. Fig. 1 shows a qualitative comparison of patch-based depth estimation models for high-resolution images. As shown, our PRO yields depth refinement results with minimal depth discontinuities along grid boundaries, whereas two recent state-of-the-art (SOTA) models still exhibit noticeable boundary artifacts. Fig. 2 illustrates a comparison of performance and efficiency between recent SOTA depth refinement models and our PRO. As observed, our PRO provides the best performance in terms of $\mathrm{D}^{3}\mathrm{R}$, AbsRel and inference speed. Our contribution is summarized as:
@@
  • We propose Grouped Patch Consistency Training (GPCT) strategy that can be well harmonized with existing DE models to mitigate the depth discontinuity problem for patch-wise DE approaches on high-resolution images. The GPCT does not require test-time ensemble, allowing a 12$\times$ faster inference speed compared to the SOTA patch-based method.
@@
  • We introduce Bias Free Masking (BFM) to selectively apply supervision signals by masking out unreliable regions in images. The unreliable regions are the regions that have inaccurately annotated depth in GT, which is often observed in the UnrealStereo4K dataset. This issue is effectively addressed by our BFM, which prevents the model from overfitting to domain-specific biases.
@@
  • Our PRO achieves SOTA performance in zero-shot evaluation on Booster [29], ETH3D [35], Middlebury 2014 [34], and NuScenes [4] across all metrics, and achieves efficient inference (12$\times$ faster than PatchRefiner with 177 patches), outperforming recent state-of-the-art depth refinement methods on high-resolution images.
@@ 2.1 Zero-shot Monocular Depth Estimation
Early MDE models [1, 25, 10, 33, 49] were trained individually on each dataset [12, 37], achieving high performance on the training data but poor generalization to unseen datasets due to the domain gap. Zero-shot MDE models that perform well on unseen images have been widely studied to improve generalization. MegaDepth [22] and DiverseDepth [48] construct large-scale datasets by collecting Internet images, improving adaptability to images from diverse scenes and conditions. They have inspired further efforts to scale up training datasets to enhance zero-shot performance. MiDaS [30] proposes a scale and shift invariant loss, which enables MDE networks to be trained on a diverse mixture of datasets, making the model robust to unseen datasets. DPT [31], Omnidata [9], and MiDaS v3.1 [2] improve DE performance by replacing CNNs with transformer architectures. DepthAnything [46] introduces a semi-supervised approach that utilizes 62M unlabeled images as pseudo labels for training. This substantial increase in data scale has led to improved generalization performance. DepthAnythingV2 [47] observes that real-world depth datasets contain label noise and lack fine details. It is trained with pseudo GT labels obtained from a teacher model trained on a synthetic dataset, achieving high-quality depth maps.
Instead of scaling up datasets, some studies utilize prior knowledge from diffusion models for DE. Marigold [19] fine-tunes a pretrained Stable Diffusion [32] model on a relatively small synthetic dataset, showing competitive results. Geowizard [11] jointly estimates depth and surface normals by leveraging self-attention across different domains to enhance geometric consistency. However, as most zero-shot MDE models are trained on low-resolution images (e.g., 384$\times$384, 518$\times$518), their performance degrades when applied to high-resolution images. When high-resolution images are downsampled to the training resolution for inference, fine details are lost. On the other hand, performing inference with a larger resolution input preserves details but degrades depth accuracy.
@@ 2.2 High-Resolution Depth Estimation
High-resolution DE models aim to predict accurate depth while preserving fine details. SMD-Net [38] utilizes an implicit function [5] to predict a mixture density for precise DE at object boundaries. Dai et al. [7] introduce a Poisson fusion-based depth map optimization method by applying guided filtering in a self-supervised learning framework. SDDR [21] proposes a self-distilled depth refinement method that generates pseudo edge labels, addressing local consistency issues and edge deformation noise through a noisy Poisson fusion process. Patch-based high-resolution DE methods select patches and merge patch-wise results to improve depth details. BoostingDepth [26] introduces a patch selection process based on edge density, and employs iterative refinement through multi-resolution depth merging. PatchFusion [23] avoids patch selection by applying a predefined partitioning scheme, and leverages shifted window self-attention to fuse global and local information. PatchRefiner [24] initializes depth estimation with a low-resolution depth map, predicts residuals, and proposes a detail and disentangling (DSD) loss for training on both real-world and synthetic datasets. However, patch-based methods inherently suffer from depth discontinuities at patch boundaries, as they emphasize local information over global consistency, or maintain depth continuity only within individual patches. To address this problem, prior works [23, 24] employ a large number of patches for test-time ensemble, which significantly slows down inference. By contrast, our PRO (Patch Refine Once) model efficiently achieves fine-grained depth refinement with only a single pass per patch during inference, despite employing a tile-based approach.
2.3 Training Data for HR Depth Estimation
High-resolution DE requires datasets with fine-grained depth details and minimal label noise. Real-world datasets [12, 37, 4, 15, 22] are often sparse due to the limitations of LiDAR or contain inaccurate depth boundaries when obtained through Structure-from-Motion (SfM). Therefore, previous research [7, 26] has trained depth networks using a small number of relatively dense real-world datasets [44, 20, 34]. Some studies [7, 21] utilize pseudo labels for self-supervision, but these labels often contain noise. Other works [23, 24] leverage synthetic datasets for more accurate depth, but suffer from poor generalization ability due to the domain gap. To take advantage of dense depth annotations in synthetic datasets without compromising generalization performance, we introduce Bias Free Masking (BFM), which selectively identifies the regions where supervision is applied. It mitigates overfitting to biases in synthetic datasets and preserves robustness in zero-shot high-resolution DE.
Figure 3: Labeling error examples in UnrealStereo4K [38] dataset.
@@ 3 Method
We describe our Patch Refine Once (PRO) model. Section 3.1 introduces the overall pipeline of PRO, followed by the two key contributions: Grouped Patch Consistency Training (GPCT) and Bias Free Masking (BFM), which are detailed in Sections 3.2 and 3.3, respectively.
@@ 3.1 Overview of Patch Refine Once (PRO)
To complement the local details, we take the following steps. First, we crop the original image $\mathbf{I}$ into $N$ grid-based patches $\{\mathbf{P}^{i}\}_{i=1}^{N}$. Each individual patch $\mathbf{P}^{i}$ is then resized and fed into $\Psi$ to estimate the corresponding fine depth map $\mathbf{D}_{\mathrm{f}}^{i}$. We extract two sets of five-level features from the decoder of $\Psi$: (i) $\mathbf{F}_{\mathrm{c}}=\{\mathbf{f}_{\mathrm{c},j}^{i}\}_{j=1}^{5}$ for the resized image, and (ii) $\mathbf{F}_{\mathrm{f}}=\{\mathbf{f}_{\mathrm{f},j}^{i}\}_{j=1}^{5}$ for the patch $\mathbf{P}^{i}$. To align $\mathbf{F}_{\mathrm{c}}$ with $\mathbf{F}_{\mathrm{f}}$, we apply the $\mathsf{ROI}$ operation [16] to extract patch-aligned features from $\mathbf{F}_{\mathrm{c}}$. The patch-level features $\mathbf{F}_{\mathrm{f}}^{i}$ and the corresponding $\mathsf{ROI}$-extracted features are subsequently passed into a fusion module within the residual prediction network $\theta$. The fusion module integrates these features using a wavelet transform [8] and shallow convolutional layers, effectively combining global and local information. See supplementary material for details of the fusion module. Finally, the refined depth map for $\mathbf{P}^{i}$, denoted as $\mathbf{D}_{\mathrm{refine}}^{i}$, is obtained as:
$\mathbf{D}_{\mathrm{refine}}^{i}=\theta(\mathsf{concat}(\mathbf{P}^{i},\mathsf{ROI}(\mathbf{D}_{\mathrm{c}},\mathbf{P}^{i}),\mathbf{D}_{\mathrm{f}}^{i})),$ (1)
where $\mathsf{ROI}(\mathbf{D}_{\mathrm{c}},\mathbf{P}^{i})$ extracts the region of interest from $\mathbf{D}_{\mathrm{c}}$ corresponding to the spatial extent of $\mathbf{P}^{i}$, yielding $\mathbf{D}_{\mathrm{c}}^{i}$. The operator $\mathsf{concat}$ denotes channel-wise concatenation. Unlike previous methods [23, 24] where separate networks are trained to predict $\mathbf{D}_{\mathrm{c}}$ and $\mathbf{D}_{\mathrm{f}}$ individually, we utilize the same pretrained zero-shot MDE network $\Psi$ and train only the residual prediction network as shown in Fig. 4-(a).